HUBBLE SPACE TELESCOPE DAILY REPORT #2280 PERIOD COVERED: 0000Z (UTC) 01/05/99 - 0000Z (UTC) 01/06/99 Daily Status Report as of 006/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Nine Sets of NIC/1/2/3 8093 (DARKs: Warming Up - Part II) The Near Infrared Camera (Cameras One, Two and Three) was used to make a series of dark calibration observations. The purpose of this proposal is to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Three Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/MA1 7471 (STIS Snapshot Search for UV-bright High Redshift Quasars) The Space Telescope Imaging Spectrograph (MA1) was used to make ultraviolet observations of the bright high redshifted quasar HS1140+350. This quasar shows no signs of strong absorption in its optical spectra, and its He-II Lambda 304 A0 line is accessible with HST. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed FOC/96 6932 (FOC Cycle 6 UV Flatfield) The Faint Object Camera (f/96) was used to make ultraviolet observations of the FFT9 region of the Orion Nebula. An internal flat and a dark calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Two Sets of NIC/2/3 8082 (NICMOS Pure Parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Three Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.9 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II parallel archive proposal continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Two Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Four Sets of NIC/1/2/3 8083 (FLATs: Warming Up - continuation) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the variable star UY-Aur as part of a T-Tauri class star survey. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2/NIC/3 7382 (The Star-Formation History of the Large Magellanic Cloud) The WF/PC-2 and the Near Infrared Camera (Camera Three) were used to make observations of the disk of the Large Magellanic Cloud (LMC-BAR-3). These data will enable us to quantitativly study the intermediate to low-mass stellar populations and the star-formation history of the LMC. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxy 3C93.1. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena; they may be used as probes of their environments; and they are unique probes of the early Universe. The observations were completed as planned, and no problems were reported. 1.15 Completed STIS/CCD/NIC/1/2/3 7566 (Demographics of Nuclear Black Holes) The Space Telescope Imaging Spectrograph (CCD) and the WF/PC-2 were used to make observations of normal galaxy NGC-4486 to try and detect a super massive black hole. A dark and wavelength calibration images were also taken. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was reset three times. A TTR was written to document a data hit at 005/2244Z whe1198 minor frames were missed. The CCS command activity concluded uneventfully at 005/2244Z. All test objectives were met with no problems. Another period of CS commanding will commence at approximately 006/1400Z. The latest analysis of NICMOS FPA temperatures indicate that the warm-up rate on Camera-3 has increased to a rate of approximately 9 degrees K per day, indicating that the thermal short has strengthened in the Camera-3 baffle. The saturation point of -78.1 degrees K is rapidly approaching and should be reached this afternoon and evening. Procedures are in place to inhibit the filter wheel commanding (probably this morning between 006/1427Z and 006/1442Z) prior to when these parameters flag out-of-limits. ESTR-2 was reconditioned at 006/0709Z per ROP RD-7. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH